OpenStudy (happykiddo):

A speeder traveling at 42 m/s passes a motorcycle policeman at rest at the side of the road. The policeman accelerates at 3.66 m/s2. To the nearest tenth of a second how long does it take the policeman to catch the speeder? Both located at the same place and time.

OpenStudy (happykiddo):

I used the formula \[S=1/2(a)(t)^2+ Vi(t)+Si\]

OpenStudy (happykiddo):

Because of the same place and time I made cop=speeder

OpenStudy (happykiddo):

So the way I did it is (1/2)(3.66m/s^2)(t^2)=(42m/s)(t)

OpenStudy (happykiddo):

I got it!! I incorrectly punched some numbers into the calculator before. But thank you for your effort : )

OpenStudy (anonymous):

The speeder is traveling at constant velocity The policeman is traveling at constant acceleration. For constant velocity I would use \[d=vt\] and for constant acceleration, I would use \[final distance=initial distance+vt+\frac{ 1 }{ 2 }at^2\] Well, what do you know? They both equals to distance traveled. So you set them equal to each other by solving for time.

OpenStudy (anonymous):

-_- lol. Right when I almost finished.

OpenStudy (happykiddo):

But still thanks!! Could you help with another problem?

OpenStudy (anonymous):

Writing in the equation sheet usually delays me.

OpenStudy (anonymous):

Yea sure.

OpenStudy (happykiddo):

I'l lpost it as a new question.

OpenStudy (anonymous):

Ok I'll be there.

OpenStudy (happykiddo):

A car traveling at 16 m/s accelerates at 3.95 m/s2 for 6 seconds. To the nearest meter how far does it travel? I used the same equation from the last question but got 77.1 meters instead of 167(correct answer).

OpenStudy (happykiddo):

\[S=\frac{ 1 }{ 2 }(a)(t)^2+Vi(t)+Si\]

OpenStudy (happykiddo):

a=3.95 t=6 Vi=0 Si=0

OpenStudy (anonymous):

No, the initial velocity is 16m/s. It does not say the vehicle started from rest.

OpenStudy (anonymous):

16m/s*6s=96m

OpenStudy (happykiddo):

Gotcha, let me try it out.

OpenStudy (happykiddo):

I got it! Thank you for the help.

OpenStudy (anonymous):

No problem!